OpenStudy (anonymous):

Samantha is going on vacation for the summer and is trying to choose between two different plans. The first plan costs $450 for 3 days at a hotel and 2 days at an amusement park. The second plan offers 5 days at the same hotel and 3 days at the amusement park for $700. The cost of 1 day at the hotel and the amusement park is the same under both plans. How much does a 1 day trip to the amusement park cost? A. $50 B. $75 C. $150 D.$200

OpenStudy (anonymous):

I'm terrible at word problems...

OpenStudy (anonymous):

x=1 day at the hotel y=1 day at the amusement park 3x+2y=450

OpenStudy (john_es):

Ok, it goes well.

OpenStudy (anonymous):

5x+3y=700

OpenStudy (john_es):

Good job ;).

OpenStudy (john_es):

Now, time for the solution. ;)

OpenStudy (anonymous):

Where I have trouble haha

OpenStudy (anonymous):

The answer is C...? :)

OpenStudy (john_es):

Try reduction again. Multiplicate the first by -3 and the second by 2, then, \[-9x-6y=-1350\\ 10x-6y=1400\] Then sum both, \[x=50\] I left the rest to you (substitute in the first equation and you'll get it) y=110

OpenStudy (john_es):

−9x−6y=−1350 10x+6y=1400 I mean

OpenStudy (anonymous):

I think I multiplied wrong...

OpenStudy (anonymous):

I either set it up wrong or messed up the multiplication

OpenStudy (john_es):

No, I missed up. The system is ok, −9x−6y=−1350 10x+6y=1400 And you obtain x=50. Then, you'll have for the y, using 450=3x+2y, 450=150+2y y=150. Again you're right, it is C.

OpenStudy (anonymous):

Multiply the first equation by 5. 15x+10y=$2250 Multiply the second by -3 -15x-9y=-$2100 Next, I added them. y=$150 :)